OpenStudy (anonymous):

1. Solve tan(x)(tan(x)-1)=0. 2. Solve on the interval [0,2pi): 2cos^2 x +3cos(x) + 1 = 0.

OpenStudy (ddcamp):

Problem 1: tan(x)(tan(x)-1) = 0 tan²(x) - tan(x) = 0 tan²(x) = tan(x) tan(x) must be 0 or 1 When does (sin(x)/cos(x)) = 0? When does it equal 1?

OpenStudy (anonymous):

Oh goodness, I don't believe I know when (sin(x)/cos(x))= 0 and 1. Can you elaborate on how to find that?

OpenStudy (ddcamp):

sin(x)/cos(x) will equal 1 when sin(x) = cos(x) sin(x)/cos(x) will equal 0 when sin(x) = 0